I am working on solve the following logarithmic equation. Round answer to 4 decimal places. The problem is 1n(x+6)+1n(x+1)=1n(x+2) I have 2 answers it ask for x={ then it will also be a decimal answer. Thanks!
Answer by edjones(8007)

1n(x+6)+1n(x+1)=1n(x+2)
e^1n(x+6)*e^1n(x+1)=e^1n(x+2)
(x+6)(x+1)=x+2
x^2+7x+6=x+2
x^2+6x+4=0
x^2+6x=-4
x^2+6x+9=9-4 completing the square
(x+3)^2=5
x+3=+-sqrt(5)
x=-3+sqrt(5)(exact answer) OR .7639 (approximate)
.
Ed